Suppose that we want to test the claim that the mean of a population is greater than 20. Also suppose that the population standard deviation is known to be 7. A sample of 49 measurements reveals a mean of 22. What is the value of the test statistic?
1
Expert's answer
2020-05-20T17:38:00-0400

"z=\frac{x−μ}{σ/\sqrt{n}}=\frac{22-20}{7/\sqrt{49}}=2.00"